Klapperich Lab:Notebook/Lab Meeting Notes/2008/09/30

From OpenWetWare
Jump to navigationJump to search
Owwnotebook icon.png Project name Report.pngMain project page
Resultset previous.pngPrevious entry      Next entryResultset next.png

30 September 2008 Lab Meeting Agenda

† No research report today.

† Announcements

  • A slow migration to 709 can now begin. There will be core equipment entering 720, we will accommodate.
  • Supply costs are out of control! Conserve.

† Flu R01

  • Mark's role outlined this week. He will make fixture for Sonali in near term. Will look at integrated chip in long term.
  • Will do RTPCR and not anti sera assays - Cathie told Chris et al. 9/29.
  • More sample back? Can we put in 1.5 ml saline, collect in 3ml VTM tube directly? Can we amend - Cathie emailed on 9/29.
  • Plaque assay status. Agar overlay problem is solved. Data later.
  • Cathie STILL has to submit the progress report to INSPIR.

† SEPSIS Project - CIMIT Bird Flu

  • Sample (raw pellets and film) from Ticona COC ordered.
  • B sub in blood is done. Lets talk about comepleting the gram positive plots so they "match."
  • PCR Pipettes bought...no one uses them for non-PCR use!!!!!!!!
  • E coli in pee/blood done.
  • SEM image analysis of pore size. Jane did preliminary analysis. Can Jane finish this up for the paper? Ongoing. Images from Hussam. Thresholding is an issue in the image processing. More planar images.

†RNA project

  • Sonali will teach Stephanie Jurkat Culture.
  • Jeff Braman will visit in early Dec.
  • We have Anirban's protocol and system etc. Issues resolved.
  • Stephanie working with Paul to make more straws.
  • SM and AC will do first experiment on Friday.


  • Mark finish control circuit (Temp, humidity, pressure) design and ordered parts
  • Fluorescence microscopy available from 720. Try beads experiment.
  • Cutter plotter ordered to make new fluidic layer
  • Jane, Cathie, Larry, Ranjith will meet soon to discuss scale up of sol gel in chip process.
  • Discuss dispensing from evap chip in a real world way.
  • Current design gives E.coli signal, clotting is still a problem
  • Change of design: funnel-shaped wells, smooth hole, make samples in pipette tips?
  • Jane working on CIMIT fellowship presentation and future plans for THESIS!

† Valve Array/Valve building (FJ)

  • Waiting for fixture still from Melissa. Hopefully this week.
  • Update: Monday 29 Sept Frank contacted Shane Conran of Superior Fluid Solutions (referred by Jess from UCSB). Will draw up preliminary design and send to us for evaluation.
  • Update: Monday 6 Oct Haven't heard anything from Shane yet. Attempting to contact again, so far without success.

† Fraunhofer: LOAC

  • Alex working with Paul. (something called personal lubricant)
  • Alex working on more toluene bonding issues.
  • Try Ticona material. (HM, JD, AG)
  • Still Still Still troubleshooting ??

† Biointerfaces group

  • Make a mold for Brett this week. When mask is made.
  • Do preliminary E. coli experiment soon.
  • Making a CR mask Tuesday afternoon.

† CIMIT- Colson Grant

  • Preliminary plan for IRB submission at Harvard in place. Target date 1/1/09
  • STAR-CD Renewal - January ($2K)


  • PCR areas defined. Protocols reviewed. Contamination issues discussed....
  • QQ tested Taqman assays. Two probe sets. One set is better than the other.
  • Contamination. Make an internal control on the chip.
  • Only 1/5 chips are good at the hot embossing step. The vaccum chuck will be checked and o-rings changed (JD)
  • Fixuture for PCR chip. The heater and the thermocouples would be on the fixture.
  • MinCheol is making Four simple PCR models to optimize temperature distribution along to the channel. Will do another model with the insulation on the bottom.
  • MinCheol is simulating DNA's Residence Time for simple PCR models. The DNA particles need to be much smaller. Also, consult the Lab on Chip paper that JD mentioned.


  • Cathie Still Working on Whitepaper almost done!
  • Send PAs to Sonali and Jaephil for sending RCA grant on to someplace else.

† Senior project

  • Megan ordering oligos and primers for the RCA. Quote. Check with Heiko on his sources.
  • Haven't seen Teddy yet.

† F31: Cochlea

  • ARO Abstract due October 1st.
  • Data summary focusing on the fundamental biological questions for next week.

† Silica Optimization (Lambda):

  • Ongoing. More concentrations. Forged chips.
  • Pressure testing on these channels when they are flowing.

hussam's update